Excel Online Freeform Activity: Hypothesis Tests for the Population Mean when o Unknown According to a recent study annual per capita consumption of milk in the United States is 22.6 gallons. Being from the Midwest, you believe milk consumption is higher there and wish to test your hypothesis. A sample of 14 individuals from the Midwestern town of Webster City was selected and then each person's milk consumption was entered into the Microsoft Excel Online file below. Use the data to set up your spreadsheet and test your hypothesis. X Open spreadsheet a. Develop a hypothesis test that can be used to determine whether the mean annual consumption in Webster City is higher than the national mean. Ho: 22.6 Ha: 22.6 b. What is a point estimate of the difference between mean annual consumption in Webster City and the national mean? 4.86 (2 decimals) c.At a = 0.01, test for a significant difference by completing the following. Calculate the value of the test statistic (2 decimals). 1.878
a. Develop a hypothesis test that can be used to determine whether the mean annual consumption in Webster City is higher than the national mean. Ho: 22.6 Ha: >22.6 b. What is a point estimate of the difference between mean annual consumption in Webster City and the national mean? (2 decimals) 4.86 c. At a 0.01, test for a significant difference by completing the following. Calculate the value of te test statistic (2 decimals) 1.878 (4 decimals) The p-value is0428 Reject the null hypothesis? No What is your conclusion? There is insufficient evidence to conclude that the population mean consu
Gallons of Milk 25.6 Anova: Single Factor 23.84 Anova: Two-Factor With R 25.25 21.2 Anova: Two-Factor Withou 17.52 19.61 Correlation 19.83 26.18 Covariance 34.97 30.2 Descriptive Statistics 28.59 20.57 Exponential Smoothing 26.94 27.24 F-Test Two-Sample for Varia
Answer #1

(a)

Correct option:

H0: 22.6

Ha: 22.6

(b)

From the given data, the following statistics are calculated:

n = 14

\bar{x} = 24.8243

s = 4.7768

So,

Point estimate of difference = 24.8243 - 22.6 = 2.22

So,

Answer is:

2.22

(c)

Test statistic is given by:

2.22 t: 4.7768/V14 = 1.74

ndf =n - 1= 14- 1 = 13

One Tail Right Side Test

By Technology,P - Value = 0.0527

Correct option:

No

Correct option:

There is insufficient evidence to to conclude that the mean population mean consumtion in Webster City is higher than national average.